OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <7, 2>, v = <21, 6>

jimthompson5910 (jim_thompson5910):

let u and v be two vectors they are orthogonal if and only if u dot v = 0 ------------------------------------------------------- they are parallel if and only if v = k*u where k is some scalar

OpenStudy (anonymous):

ok so how do i find u dot v

jimthompson5910 (jim_thompson5910):

u = <a,b> v = <c,d> u dot v = a*c + b*d the * means multiplication between two scalars u,v are vectors a,b,c,d are scalars

jimthompson5910 (jim_thompson5910):

example u = <2,3> v = <7,9> u dot v = 2*7 + 3*9 = 14 + 27 = 41

jimthompson5910 (jim_thompson5910):

"dot" refers to the dot product of two vectors

OpenStudy (anonymous):

Ok so for my problem u dot v = 159.

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

since that result isn't 0, the vectors aren't orthogonal

OpenStudy (anonymous):

How would I determine if they are parallel or not?

jimthompson5910 (jim_thompson5910):

can you find some scalar k that makes v = k*u true?

jimthompson5910 (jim_thompson5910):

does k = 5 make v = k*u true? ie is v = 5*u true?

jimthompson5910 (jim_thompson5910):

Alternate method Let u = <a,b> v = <c,d> u and v are parallel if and only if a/c = b/d is true

OpenStudy (anonymous):

Okay I got 7/ 21 = 2/6 so they are parallel, thanks!

jimthompson5910 (jim_thompson5910):

yep they are parallel